a van leaves a store at 2:00 going 35 miles an hour
a car leaves the same store at 2:45 going 50 miles an hour
how long will it take for the car to pass the van

Answer by ankor@dixie-net.com(12701)

a van leaves a store at 2:00 going 35 miles an hour
a car leaves the same store at 2:45 going 50 miles an hour
how long will it take for the car to pass the van
:
Since we are dealing in hrs here, change 45 min: 45/60 = .75 hrs
:
Let t = time traveled for car to pass the van
then
(t+.75) = time traveled by the van when this happens
:
We know that when the car passes the van, they will have traveled the same distance.
We can make a simple distance equation from the this fact.
:
Distance = speed * time
:
Car dist = van dist
50t = 35(t+.75)
:
50t = 35t + 26.25
:
50t - 35t = 26.25
:
15t = 26.25
:
t = 26.25/15
:
t = 1.75 hrs, or 1 hr 45 min, for the car to pass the van
:
:
WE can check our solution by seeing if both vehicles have traveled the same distance.
50(1.75) = 87.5 mi
35(1.75+.75) =
35(2.5) = 87.5 also confirms our solution
:
How about this? Did it make sense to you?